# Rationale-based Ensemble of Multiple QA Strategies for Zero-shot Knowledge-based VQA

### v2

@@ Abstract
Knowledge-based Visual Qustion-answering (K-VQA) necessitates the use of background knowledge beyond what is depicted in the image. Current zero-shot K-VQA methods usually translate an image to a single type of textual decision context and use a text-based model to answer the question based on it, which conflicts with the fact that K-VQA questions often require the combination of multiple question-answering strategies. In light of this, we propose Rationale-based Ensemble of Answer Context Tactics (REACT) to achieve a dynamic ensemble of multiple question-answering tactics, comprising Answer Candidate Generation (ACG) and Rationale-based Strategy Fusion (RSF). In ACG, we generate three distinctive decision contexts to provide different strategies for each question, resulting in the generation of three answer candidates. RSF generates automatic and mechanistic rationales from decision contexts for each candidate, allowing the model to select the correct answer from all candidates. We conduct comprehensive experiments on the OK-VQA and A-OKVQA datasets, and our method significantly outperforms state-of-the-art LLM-based baselines on all datasets.
Rationale-based Ensemble of Multiple QA Strategies for Zero-shot Knowledge-based VQA
Miaoyu Li, Haoxin Li, Zilin Du, and Boyang Li College of Computing and Data Science, Nanyang Technological University, Singapore
@@ 1 Introduction
An effective technique for zero-shot VQA is to translate an image to a textual decision context for a text-based model, such as an LLM, which answers the question using the generated decision context (Yang et al., 2022; Tiong et al., 2022). For example, Yang et al. (2022) translate the image to captions; Guo et al. (2023) further generate question-answer pairs based on captions. To cater to the knowledge-intensive nature of K-VQA, Cao and Jiang (2024) generate a paragraph describing relevant background knowledge, from which an LLM extracts the answer. Such approaches rely on a single question-answering strategy using a single type of decision context.
However, we observe that K-VQA datasets often require the joint application of multiple strategies. For some questions, merely the captions are sufficient for answering, whereas other questions may require background knowledge in either short or long forms. As an illustration, in Fig 1 we show three questions and three AI-generated decision contexts for each question, including captions, short-form knowledge of a single sentence, and long-form knowledge containing multiple sentences. In each question, only one decision context leads to the right answer.
@@
In light of this, we propose Rationale-based Ensemble of Answer Context Tactics (REACT) to achieve dynamic ensemble of multiple question-answering tactics. REACT comprises an Answer Candidate Generation (ACG) phase and a Rationale-based Strategy Fusion (RSF) phase. Both phases use an in-context-learning LLM without any training. ACG generates three decision contexts for each K-VQA question, including image captions, short-form knowledge, and long-form knowledge, with increasing amount of background knowledge. The ACG LLM then generates three answer candidates, one from each decision context.
The RSF LLM works as an arbitrator that selects one correct answer from the candidates provided by ACG. In order for the RSF LLM to make informed choices, we provide two complementary types of rationales for each answer candidate. The first type, automatic rationale, is generated by asking the LLM to summarize the reasoning behind the answer in one sentence. The second type, mechanistic rationale, is a sentence from the decision context that has the most contribution to the answer, as identified by a model interpretation technique, GradCam (Selvaraju et al., 2017). These brief rationales preserve the most relevant information in the original decision contexts, and prevent irrelevant information from misleading the RSF LLM. The RSF LLM receives two rationales for each answer candidate and selects one final answer from the three candidates, achieving dynamic question-answering strategy ensemble.
On two popular K-VQA datasets, OKVQA and A-OKVQA, REACT outperforms state-of-the-art baselines that use frozen LLMs without training by 2.6% to 4.7%. Ablation studies reveal that the fusion of three answers improves performance over the best single answer strategy by 1.1% on OK-VQA and 1.6% on A-OKVQA. Further, using both the automatic and mechanistic rationales obtains gains of 1.2% and 1.4%, respectively.
@@
  • We identify the need for combined use of multiple question-answering strategies in K-VQA, and propose REACT, which fuses multiple answers and answer strategies dynamically using frozen in-context learning LLMs, achieving state-of-the-art performance on zero-shot OKVQA and A-OKVQA.
@@ 2 Related Work
Another type of zero-shot method converts visual information to textual descriptions and applies frozen LLMs to answer questions based on these descriptions alone. We classify these methods based on the number of types of descriptions used and whether answer ensemble is employed. Most methods (Yang et al., 2022; Tiong et al., 2022; Guo et al., 2023; Chen et al., 2024) translate images into a single type of description: captions, and generate one answer from them. Chen et al. (2024) use two different LVLMs to generate different captions and answer candidates, and one LLM to choose the final answer. To accommodate the need for external knowledge in K-VQA, Cao and Jiang (2024) generate two types of descriptions, captions and long-form knowledge, for each question. A single answer is obtained based on the concatenated caption-knowledge context. In contrast, our method focuses on the explicit answer selection process from multiple answer candidates.
Figure 2: The framework of REACT. The ACG module contains three question-answering strategies, which generate three different decision contexts for the answer-predicting LLM. The short-form knowledge contains a single sentence, whereas the long-form knowledge consists of a paragraph of background information. The caption-only strategy generates only image captions and no background knowledge. For each strategy, the LLM generates an answer candidate. In the RSF module, for each candate, we generate an Automatic (A.) Rationale and a Mechanistic (M.) Rationale, which guide the LLM in selecting the best answer.
@@ 3 Method
The goal of our method is to jointly utilize multiple decision contexts to answer each question and achieve a dynamic ensemble of these different strategies. To accomplish this, we propose two modules: Answer Candidate Generation (ACG) and Rationale-based Strategy Fusion (RSF). This section provides detailed descriptions of them.
3.1 Answer Candidate Generation
In ACG, we generate three types of decision contexts for each question, including captions, short-form knowledge consisting of one sentence, and long-form knowledge containing multiple sentences. They provide three different strategies for each question, resulting in three answer candidates.
@@ 3.1 Answer Candidate Generation
Since LLMs can only process textual information, we use an off-the-shelf image captioning model to transform an image into multiple textual captions. To ensure the relevance to the question, we employ the question-guided caption generation mechanism in PNP-VQA (Tiong et al., 2022), which first identifies the image patches most relevant to the question and then generates captions only from these relevant patches. The prompt for the caption-only strategy is shown in Tab. 4 of Appendix A.
To achieve in-context learning, we follow Img2LLM (Guo et al., 2023) and generate Question-answer (QA) pairs from the image captions several. The QA pairs serve as demonstration of the QA task for LLM in-context learning. As the task demonstration is necessary for LLMs to perform QA, we include these QA pairs in the prompt of all three strategies of ACG.
@@
Short-form knowledge is highly relevant to the visual question, making it accurate. However, due to its short length, it may not be comprehensive enough to get the correct answer.
@@
In this strategy, we remove the length constraint “in a single sentence” from the prompt, so that the LLM can generate as much text as it wants. In most cases, this leads to one or two paragraphs, which ideally would capture comprehensive background knowledge related to the question. The prompts for generating long-form knowledge and utilizing it to generate answer are detailed in Tab. 5 and Tab. 4 of Appendix A.
The short-form knowledge and the long-form strategy are synergistic. Constrained to a single sentence, the short-form knowledge context is succinct but may miss some relevant information. In comparison, the long-form knowledge strategy prioritizes knowledge recall over precision, and may contain many irrelevant facts that potentially distracts the LLM. Their synergy is further demonstrated in the experiments (§4.3 and Fig. 4). Balancing different strategies is hence important for high performance.
3.2 Rationale-based Strategy Fusion
Since not all question-answering strategies are suitable for a given question, selecting the correct answer is crucial. To allow the LLM to make informed choices, RSF generates two types of rationales, automatic and mechanistic rationales, for each answer candidate. These brief rationales retain only the most relevant information from the decision context, serving as the basis for answer selection, thus mitigating the risk of misleading the LLM with a plethora of irrelevant information.
@@ 3.2 Rationale-based Strategy Fusion
To generate the automatic rationale, we feed the original question, the decision context, and the predicted answer to an LLM and ask the LLM to summarize the rationale behind the answer to a single sentence. The LLM is not restricted to selecting a sentence from the decision context and performs open-ended generation. The prompt for generating automatic rationales is shown in Tab. 6 of Appendix A.
@@
Though the automatic rationale is easy to acquire and often reasonable, there are occasions when the LLM generate incorrect rationales by ignoring part of the decision context or hallucinating. Although this could be partially alleviated by exhaustive prompt engineering, we opt for a more systematic solution, which is to obtain the rationale through a mechanistic interpretation of the answer-generating LLM.
Inspired by GradCAM (Selvaraju et al., 2017), we devise a method to compute the contribution of each sentence in the decision context to the answer candidate. We first introduce some notations. The prompt to the LLM, including the instruction and the decision context (captions or knowledge statements) contains $N_{P}$ tokens. The answer generated by the LLM contains $N_{A}$ tokens. When generating the $k$-th answer token $w_{k}$, the Transformer-based LLM attends to all previous $N_{k}=N_{P}+k-1$ tokens. We extract the attention weight vector $\bm{a}^{(h)}\in\mathbb{R}^{N_{k}}$ from the $h$-th attention head. The components of $\bm{a}^{(h)}$ sum up to 1, $\sum_{i=1}^{N_{k}}a^{(h)}_{i}=1$.
We seek the contribution of every token preceding the $k$-th answer token. Instead of the raw attention scores, which may be redundant and inaccurate, we weigh the attention scores with its gradient from the probability $p_{k}$ of the predicted answer token $w_{k}$. We disregard negative gradients as we focus on positive contributions. Formally, the relevance scores of each of the $N_{k}$ token is
@@
Question-answering Strategy Fusion.
In the final fusion stage, we provide all three answer candidates with corresponding rationales, including both the automatic and the mechanistic rationales, in the prompt. The LLM is instructed to select the best answer, thereby achieving dynamic strategy ensemble. To ensure the model perceives the image content during fusion, we also include caption-generated QA pairs in the prompt. The exact prompt for question-answering strategy fusion is in Tab. 7 of Appendix A.
@@ 4 Experiment
KGenVQA UnifiedQA 11B $\times$ long 45.4 39.1 -
@@
[…]
REACT Gemma 7B long & short 47.6 47.3 46.8
REACT Mistral 7B long & short 49.2 47.5 48.6
@@ 4.1 Setup
We evaluate our method on two mainstream K-VQA datasets: OK-VQA (Marino et al., 2019) and A-OKVQA (Schwenk et al., 2022). Questions in both datasets require knowledge beyond the images to answer. We utilize the test split of OK-VQA and the val split and test split of A-OKVQA for evaluation. These splits contain 5,046, 1,100, and 6,700 questions, respectively. We follow the official evaluation protocols of direct answer and report VQA scores for each dataset.
@@
We compare REACT with previous zero-shot K-VQA methods without training. The methods using frozen LLMs can be roughly divided into two categories: (1) Methods employing the caption-only strategy, PICa (Yang et al., 2022), PNP-VQA (Tiong et al., 2022), Cola-Zero (Chen et al., 2024), Img2LLM (Guo et al., 2023), and RQprompt (Lan et al., 2023). (2) Methods employing the long-form knowledge strategy. The only method in this category is KGenVQA (Cao and Jiang, 2024). Unlike translating visual information into textual descriptions for frozen LLMs, the concurrent method REPARE (Prasad et al., 2024) directly utilizes an LVLM as QA model [^1] The REPARE+LLaVA-1.5 model does not meet the zero-shot requirement, as LLaVA-1.5 is pretrained on OK-VQA and A-OKVQA. Hence, we omit that result from the comparison..
@@ 4.2 Main Results
We observe that REACT establishes a new state-of-the-art on LLM-based zero-shot KVQA. Compared to previous best scores achieved by frozen LLMs, REACT with Mistral-7B outperforms them with tremendous margins of 2.8%, 2.6% and 4.7% on OK-VQA, A-OKVQA val split and test split, respectively. These results strongly demonstrate the effectiveness of our approach.
We also compare against a concurrent method, REPARE (Prasad et al., 2024), which directly prompts an end-to-end trained LVLM. Since REPARE does not translate image to text, it avoids any information loss in the process. Nevertheless, REACT remains highly competitive, outperforming all results of REPARE but that from MiniGPT-4-13B with a much larger LLM, and the difference is merely 0.4%. The fact that a blind LLM, when prompted properly, is competitive with an well-prompted LVLM at a visual task is surprising to us. As the text-as-visual-representation approach can benefit from advances in LLMs without extensive vision-language alignment training, REACT offers strong practical benefits.
@@
Table 2: Ablation study results of ACG with Mistral-7B. Cap. denotes captions. SK and LK are short-form knowledge and long-form knowledge, respectively.
Figure 4: The success rate of the other two strategies over questions successfully answer by each strategy. In left group, we show the percentage of correctly answered questions by SK and LK out of all the questions that are correctly answered by CO. The other two bar groups are drawn similarly.
4.3 Ablation Studies
Analysis of ACG.
Comparing #1 with #2 and #3, we observe that the short-form knowledge strategy and long-form knowledge strategy achieve better performance than caption-only strategy, demonstrating that generated background knowledge is advantageous in K-VQA. After fusing caption-only and short-form knowledge in #4, the performance on OKVQA and A-OKVQA improves by 2.1% and 2.8% over #1. This demonstrates the importance of strategy ensemble and the effectiveness of the short-form knowledge strategy. The same phenomenon can be observed in #1 and #5. Finally, the full REACT system, #6, achieves the best performance.
Further, we quantify how much each QA strategy complements others. In Fig. 4, for each QA strategy, we first extract all questions that this strategy can successfully answer by itself. Out of these questions, we show the success rate of the other two strategies. The results indicate that out of the questions answerable by one strategy, 15-19% cannot be answered by another strategy. Hence, each QA strategy is indispensable for the success of REACT.
@@ 4.3 Ablation Studies
Table 3: Ablation study of RSF with Mistral-7B. MR and AR denote mechanistic rationale and automatic rationale, respectively.

### v4

# Diversify, Rationalize, and Combine: Ensembling Multiple QA Strategies for Zero-shot Knowledge-based VQA

@@ Abstract
Knowledge-based Visual Qustion-answering (K-VQA) often requires the use of background knowledge beyond the image. However, we discover that a single knowledge generation strategy is often insufficient for all K-VQA questions. To this end, we propose Diversification, Evidence Truncation, and Combination for Knowledge-based Elucidation (DietCoke), which utilizes a bundle of complementary question-answering tactics and aggregates their answers using textual rationales. DietCoke comprises of three stages: diversification, rationalization, and ensemble. The diversification stage generates three distinctive decision contexts, each leading to its own answer candidate. The rationalization stage generates two rationales, the automatic rationale and the mechanistic rationale, for each answer candidate using decorrelated techniques. Finally, in the ensemble stage, an LLM informed by the rationales selects one answer from the three candidates. Experiments show that DietCoke significantly outperforms state-of-the-art LLM-based baselines by 2.8% on OK-VQA and 4.7% on A-OKVQA and that the strategies in the ensembles are highly complementary. Code is available at: https://github.com/limiaoyu/DietCoke
Diversify, Rationalize, and Combine: Ensembling Multiple QA Strategies for Zero-shot Knowledge-based VQA
Miaoyu Li, Haoxin Li, Zilin Du, and Boyang Li College of Computing and Data Science, Nanyang Technological University, Singapore {miaoyu.li, haoxin003, zilin003, boyang.li}@ntu.edu.sg
@@ 1 Introduction
An effective technique for zero-shot VQA is to translate an image to a textual decision context for a text-based model, such as an LLM, which the answers the question using the context (Yang et al., 2022; Tiong et al., 2022). For example, Yang et al. (2022) translate the image to captions; Guo et al. (2023) further generate question-answer pairs to demonstrate the VQA task. To cater to the knowledge-intensive nature of K-VQA, Cao and Jiang (2024) generate a paragraph describing relevant background knowledge, from which an LLM extracts the answer. These approaches rely on a single question-answering strategy using a single type of decision context.
Interestingly, we observe that one LLM strategy is often not sufficient for K-VQA datasets. For some questions, merely the captions are sufficient for finding the right answer. Other questions benefit from knowledge extracted from LLMs, but it is often not clear if a short, one-sentence knowledge statement or a paragraph would beget the right answer from the LLM. As an illustration, in Fig. 1 we show three questions and three AI-generated decision contexts for each question, including captions, short-form knowledge as a single sentence, and long-form knowledge, which contains multiple sentences. In each question, only one decision context leads to the right answer.
@@
In light of this, we propose Diversification, Evidence Truncation, and Combination for Knowledge-based Elucidation (DietCoke) to achieve dynamic ensemble of different question-answering strategies. DietCoke comprises a diversification phase, a rationalization phase, and an ensemble phase. In each phase, we use an in-context-learning LLM without training. In diversification, we generate three decision contexts for each K-VQA question, including image captions, short-form knowledge, and long-form knowledge, with an increasing amount of background knowledge. The LLM generates three answer candidates, one from each decision context.
In the ensemble phase, the LLM selects one correct answer from the three candidates. To encourage informed decisions, we design the rationalization phase, which generates two complementary types of rationales for each answer candidate. A rationale summarizes the portion of a decision context that can support the correctness of the answer. The first type, automatic rationale, is generated by simply asking an LLM to summarize the reasoning behind the answer into one sentence. The second type, mechanistic rationale, is one sentence from the decision context that contributes the most to the generation of the answer, as identified by a model interpretation technique, GradCam (Selvaraju et al., 2017). These brief rationales preserve the most relevant information in the original decision contexts, and prevent irrelevant information from misleading the ensemble LLM. Informed by the rationales, the LLM in the ensemble phase selects one final answer from the candidates, achieving dynamic ensemble of the question-answering strategies.
On two popular K-VQA datasets, OKVQA and A-OKVQA, DietCoke outperforms state-of-the-art baselines that use frozen LLMs without training by 2.6% to 4.7%. Ablation studies reveal that the fusion of three answers improves performance over the best single answer strategy by 1.1% on OK-VQA and 1.6% on A-OKVQA. Further, adding both the automatic and mechanistic rationales obtains gains of 1.2% and 1.4%, respectively on the two datasets.
The success of DietCoke can be understood from multiple perspectives. First, DietCoke creates an effective ensemble by employing decorrelated QA strategies and rationalization strategies. Classic theory on ensemble learning Clemen and Winkler (1985); Breiman (2001) indicates that an ensemble of weak classifiers becomes more powerful as the individual classifiers become less correlated and more diverse. Intuitively, different QA contexts lead to diverse answers. Also, disparate mechanisms used by the two rationalization strategies utilize reduce correlation. In §4.3, we perform a series of ablation studies that demonstrate the synergy among the three QA strategies and the two rationalization strategies.
Second, rationales in DietCoke can be understood from a chain-of-thought (CoT) perspective. Recently, Li et al. (2024) prove that chain-of-thought prompting expands the circuit complexity of problems solvable by an LLM from the class $AC^{0}$ with constant circuit depth to the class $P/poly$ with polynomial circuit depth. This is made possible by saving intermediate computational results as CoT tokens. Analogically, providing the QA contexts (i.e., intermediate results) to the answer-selecting ensemble LLM should be beneficial. However, in practice we observe that long decision contexts may contain many irrelevant facts that can mislead the ensemble LLM. By summarizing the contexts into rationales, we provide abridged chains-of-thoughts that expand circuit depth while avoiding misleading tokens.
@@
  • We identify the need for combined use of multiple question-answering strategies in K-VQA, and propose DietCoke, which fuses multiple answers and answer strategies dynamically using frozen in-context learning LLMs, achieving state-of-the-art performance on zero-shot OKVQA and A-OKVQA.
@@ 2 Related Work
Another type of zero-shot method converts visual information to textual descriptions and applies frozen LLMs to answer questions based on these descriptions alone. We classify these methods based on the number of types of descriptions used and whether answer ensemble is employed. Most methods (Yang et al., 2022; Tiong et al., 2022; Guo et al., 2023; Chen et al., 2024) translate images into a single type of description: captions, and generate one answer from them. Chen et al. (2024) uses two different LVLMs to generate different captions and answer candidates, and one LLM to choose the final answer. To accommodate the need for external knowledge in K-VQA, Cao and Jiang (2024) generate two types of descriptions, captions and long-form knowledge and concatenates into one prompt. Hence, we consider it to be a single-strategy method. In contrast, DietCoke explicitly combines different QA strategies by selecting from their answers.
Figure 2: The framework of DietCoke. The diversification phase contains three question-answering strategies, which generate three different decision contexts for the answer-predicting LLM. The short-form knowledge contains a single sentence, whereas the long-form knowledge consists of a paragraph of background information. The caption-only strategy generates only image captions and no background knowledge. For each strategy, the LLM generates an answer candidate. In the rationalization phase, for each candidate, we generate an Automatic (A.) Rationale and a Mechanistic (M.) Rationale, which guide the ensemble LLM in selecting the best answer.
@@ 3 Method
The goal of our method is to jointly utilize multiple decision contexts to answer each question and achieve a dynamic ensemble of these different strategies. To accomplish this, we propose DietCoke, which consists of three phases: diversification, rationalization and ensemble. This section provides detailed descriptions of them.
3.1 Diversification
In diversification phase, we generate three types of decision contexts for each question, including captions, short-form knowledge consisting of one sentence, and long-form knowledge containing multiple sentences. They provide three different strategies for each question, resulting in three answer candidates.
@@ 3.1 Diversification
Since LLMs can only process textual information, we use an off-the-shelf image captioning model to transform an image into multiple textual captions. To ensure the relevance to the question, we employ the question-guided caption generation mechanism in PNP-VQA (Tiong et al., 2022), which first identifies the image patches most relevant to the question and then generates captions only from these relevant patches. The prompt for the caption-only strategy is shown in Tab. 4 of Appendix B.
To achieve in-context learning, we follow Img2LLM (Guo et al., 2023) and generate Question-answer (QA) pairs from the image captions. The QA pairs serve as demonstration of the QA task for LLM in-context learning. As the task demonstration is necessary for LLMs to perform QA, we include these QA pairs in the prompt of all three strategies of diversification.
@@
We observe that short-form knowledge tends to be highly relevant to the visual question. However, due to its short length, it may not be comprehensive enough to get the correct answer.
@@
In this strategy, we remove the length constraint “in a single sentence” from the prompt, so that the LLM can generate as much text as it wants. In most cases, this leads to one or two paragraphs, which ideally could capture comprehensive background knowledge related to the question. The prompts for generating long-form knowledge and utilizing it to generate answer are detailed in Tab. 5 and Tab. 4 of Appendix B.
The short-form knowledge and the long-form knowledge are synergistic. Constrained to a single sentence, the short-form knowledge context is succinct but may leave out relevant information. In comparison, the long-form knowledge strategy prioritizes knowledge recall over precision, and may contain many irrelevant facts that potentially distract the LLM. Their synergy is further demonstrated in the experiments (§4.3 and Fig. 4). Balancing different strategies is hence important for high performance.
3.2 Rationalization
Since not all question-answering strategies are suitable for a given question, selecting the correct answer is crucial. To allow the LLM to make informed choices, the rationalization phase generates two types of rationales with different methods for each answer candidate, which we call automatic rationales and mechanistic rationales. These brief rationales retain only the most relevant information from the decision context, thus mitigating the risk of the LLM being misled by a plethora of irrelevant information.
@@ 3.2 Rationalization
To generate the automatic rationale, we feed the original question, the decision context, and the predicted answer to an LLM and directly ask the LLM to summarize the rationale behind the answer to a single sentence. The LLM is not restricted to selecting a sentence from the decision context and performs open-ended generation. The prompt for generating automatic rationales is shown in Tab. 6 of Appendix B.
@@
Though the automatic rationale is easy to acquire and often reasonable, there are occasions when the LLM generate incorrect rationales by ignoring part of the decision context or hallucinating (see examples in Fig. 8). Although this could be partially alleviated by exhaustive prompt engineering, we opt for a more systematic solution, which is to obtain the rationale through mechanistic interpretation of the answer-generating LLM.
Inspired by GradCAM (Selvaraju et al., 2017), we devise a method to compute the contribution of each sentence in the decision context to the answer candidate. We first introduce some notations. The prompt to the LLM, including the instruction and the decision context (captions or knowledge statements), contains $N_{P}$ tokens. The answer generated by the LLM contains $N_{A}$ tokens. When generating the $k$-th answer token $w_{k}$, the Transformer-based LLM attends to all previous $N_{k}=N_{P}+k-1$ tokens. We extract the attention weight vector $\bm{a}^{(h)}\in\mathbb{R}^{N_{k}}$ from the $h$-th attention head. The components of $\bm{a}^{(h)}$ sum up to 1, $\sum_{i=1}^{N_{k}}a^{(h)}_{i}=1$.
We seek the contribution of every token preceding the $k$-th answer token. Instead of the raw attention scores, which may be redundant and inaccurate, we weigh the attention scores with its gradient from the probability $p_{k}$ of the predicted answer token $w_{k}$. We disregard negative gradients as we focus on positive contributions. Formally, the relevance scores of each of the $N_{k}$ tokens is
@@
3.3 Ensemble
In the final ensemble stage, we provide all three answer candidates with corresponding rationales, including both the automatic and the mechanistic rationales, in the prompt. The LLM is instructed to select the best answer, thereby achieving dynamic strategy ensemble. To ensure the model perceives the image content during ensemble, we also include caption-generated QA pairs in the prompt. The exact prompt for question-answering strategy fusion is in Tab. 7 of Appendix B.
@@ 4 Experiment
KGenVQA UnifiedQA 11B long 45.4 39.1 -
@@
DietCoke Gemma 7B long & short 47.6 47.3 46.8
DietCoke Mistral 7B long & short 49.2 47.5 48.6
@@
REPARE MiniGPT-4 13B - - - 47.9 -
@@ 4.1 Setup
We evaluate our method on two mainstream K-VQA datasets: OK-VQA (Marino et al., 2019) and A-OKVQA (Schwenk et al., 2022). Questions in both datasets require knowledge beyond the images to answer. We utilize the test split of OK-VQA and the validation split and the test split of A-OKVQA for evaluation. These splits contain 5,046, 1,100, and 6,700 questions, respectively. We follow the official evaluation protocols of direct answer and report VQA scores for each dataset.
@@
We compare DietCoke with previous zero-shot K-VQA methods without training. The methods using frozen LLMs can be roughly divided into two categories: (1) Methods employing the caption-only strategy, PICa (Yang et al., 2022), PNP-VQA (Tiong et al., 2022), Cola-Zero (Chen et al., 2024), Img2LLM (Guo et al., 2023), and RQprompt (Lan et al., 2023). (2) Methods employing the long-form knowledge strategy. The only method in this category is KGenVQA (Cao and Jiang, 2024). Unlike translating visual information into textual descriptions for frozen LLMs, the concurrent method REPARE (Prasad et al., 2024) directly utilizes an LVLM as QA model [^1] The REPARE+LLaVA-1.5 model does not meet the zero-shot requirement, as LLaVA-1.5 is pretrained on OK-VQA and A-OKVQA. Hence, we omit that result from the comparison..
@@ 4.2 Main Results
DietCoke establishes a new state-of-the-art on LLM-based zero-shot KVQA. Compared to previous best scores achieved by frozen LLMs, DietCoke with Mistral-7B outperforms them with large margins of 2.8%, 2.6% and 4.7% on the OK-VQA, A-OKVQA validation split and test split, respectively. These results strongly demonstrate the effectiveness of our approach.
We also compare against a concurrent method, REPARE (Prasad et al., 2024), which directly prompts an end-to-end trained LVLM. Since REPARE does not translate image to text, it avoids any information loss in the process. Nevertheless, DietCoke remains highly competitive, outperforming all results of REPARE but that from MiniGPT-4-13B with a much larger LLM, and the difference is merely 0.4%. The fact that a blind LLM, when prompted properly, is competitive with a well-prompted LVLM at a visual task is surprising to us. As the text-as-visual-representation approach can benefit from advances in LLMs without extensive vision-language alignment training, DietCoke offers strong practical benefits.
We also compare the time cost of DietCoke with Img2LLM on Mistral-7B. We randomly select 50 questions from the OKVQA test split and 50 from the A-OKVQA val split. We report the average time spent on each question using two RTX 3090 24G GPUs. Img2LLM takes 1.7 seconds. If we maximize parallelism of LLM calls, DietCoke takes 15.8 seconds. In completely sequential execution, DietCoke takes 29.5 seconds. DietCoke provides another point on the time-accuracy Pareto front and allows a VQA system to trade compute for answer quality when additional compute is available.
@@
Table 2: Ablation study results of diversification with Mistral-7B. Cap. denotes captions. SK and LK are short-form knowledge and long-form knowledge, respectively.
Figure 4: The success rate of the other two strategies on all questions successfully answered by one strategy. In left group, we show the percentage of correctly answered questions by SK and LK out of all the questions that are correctly answered by CO. The other two bar groups are drawn similarly.
@@
Table 3: Ablation study of rationalization with Mistral-7B. MR and AR denote mechanistic rationale and automatic rationale, respectively.
4.3 Ablation Studies
The QA Strategies.
Comparing #1 with #2 and #3, we observe that the short-form knowledge strategy and long-form knowledge strategy achieve better performance than caption-only strategy, demonstrating that generated background knowledge is advantageous in K-VQA. After fusing caption-only and short-form knowledge in #4, the performance on OKVQA and A-OKVQA improves by 2.1% and 2.8% over #1. This demonstrates the importance of strategy ensemble and the effectiveness of the short-form knowledge strategy. The same phenomenon can be observed in #1 and #5. Finally, the full DietCoke system, #6, achieves the best performance.
Further, we quantify how much each QA strategy complements others. In Fig. 4, for each QA strategy, we first extract all questions that this strategy can successfully answer by itself. Out of these questions, we show the success rate of the other two strategies. The results indicate that out of the questions answerable by one strategy, 15-19% cannot be answered by another strategy. Hence, each QA strategy is indispensable for the success of DietCoke.